doglover12:

The simplified form of the expression (5.23x + 3.76) − (3.67x − 6.39) is .

COLOJEUMEDY:

1.56x+10.15

justus:

Use distributive property and combine like terms.

4evablood:

1.56x+10.15 distribute the subtraction sign, and add the like terms.

supie:

\(5.23x+3.76−(3.67x−6.39)\) Distribute the ` − (subtraction) ` sign : \(5.23x+3.76+−1(3.67x−6.39)=\) \(5.23x+3.76+−1(3.67x)+(−1)(−6.39)=\) \(5.23x+3.76+−3.67x+6.39\) Combine Like Terms: \(5.23x+3.76+−3.67x+6.39=\) \((5.23x+−3.67x)+(3.76+6.39)=\) \(1.56x+10.15\)